The sum of Peter and Paul ages is 24 six years ago Peter was three times as old as Paul. How old Paul will be in five years from now?
Answer by ikleyn(52775)

.
Six years ago, when the sum of Peter and Paul ages was 24 - 2*6 = 12,  Peter was three times as old as Paul.


It means that Paul was 3 years, while Peter was 9 years old at that time.


Hence, Paul is 3+6 = 9 years old now, while Peter's age is 9+6 = 15 years today.


In 5 years, Paul will be 9+5 = 14, while Peter will be 15+5 = 20 years old.    ANSWER.

Solved   //   Mentally,  without using equations.

-------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.